Euromold specializes in manufacturing molded a plastic panel to be fitted on car doors. The blueprint specification for the thickness of a high-demand model calls for 0.20625 ± 0.00275 inch. It costs an estimated $180 to scrap a part that is outside of specification. The thickness measure for the unit just completed is 0.20823.

Required:

1. Calculate the value of k, the cost coefficient in the Taguchi Quality Loss Function (QLF) for the above situation (rounded to nearest whole number).

2. Calculate the estimated amount of quality cost (or loss) for the unit in question, L(x) (where x = 0.20823 inches); round final answer to two (2) decimal places.
Answer #1

Answer :

(1). The Taguchi loss Function for Euromold specializes in manufacturing is

L(x) = k(x- T)2

$180 = k(0.00275)2

k = 23801.65

L(x) = k(x - T)2 = 23801.65 (x- T)2

(2).

L(x) = k (x - T)2 = 23801.65 (x - T)2

Where x = 0.20823

L(0.20823) = 23801.65 (0.20823)2 = $1032.03